### v3

# Transition Metal Dichalcogenides Multijunction Solar Cells Toward the Multicolor Limit

Abstract
Transition metal dichalcogenides (TMDs) and other van der Waals (vdW) semiconductors enable transfer-printed, lattice–mismatch-free stacking of many photovoltaic junctions, motivating a re-examination of multijunction detailed-balance limits under realistic material and optical constraints. Here, we develop a unified thermodynamic framework for a multijunction photovoltaic device, which can define a clear set of device-window constraints, optical boundary conditions, and luminescence/entropy penalties and therefore define how closely any realistic multijunction photovoltaic device can approach multicolor limit. By applying it to a conservative TMD bandgap window (1.0–2.1 eV), we show that the accessible bandgap window imposes a large-junction number ($N$) efficiency limit: under full concentration, unconstrained ladders approach 84.5% at $N=50$, whereas the TMD window plateaus near 63.4%. This efficiency plateau is set by photons outside the bandgaps, so radiative quality and optics dominate beyond $N=5$ junctions for realistic transfer-printed device stacks. We identify an experimentally achievable $N=5$ ladder $E_{g}\approx(2.10,1.78,1.50,1.24,1.00)$ eV and map each rung to candidate vdW/TMD absorbers. Using reciprocity and luminescence thermodynamics, we quantify penalties from finite external radiative efficiency, two-sided emission, and luminescent coupling, and introduce the upward-emitted luminescence power as an indicator of entropy-loss proxy. Incorporating excitonic absorptance and nanophotonic thickness bounds yields practical thickness and light-management targets for transfer-printed stacks. Finally, inserting an idealized nonreciprocal multijunction model into the reciprocity-optimized ladders provides conservative efficiency advantage estimates, which are consistent with negligible benefit for single junctions but measurable efficiency gains for multijunction TMD devices.
keywords:
multijunction photovoltaics , transition metal dichalcogenides , detailed balance , external radiative efficiency , luminescent coupling
1 Introduction
The Shockley–Queisser (SQ) detailed-balance limit for a single-junction solar cell formalizes the thermodynamic tradeoffs between (i) transmission loss from sub-bandgap photons, (ii) thermalization loss from above-bandgap photons, and (iii) radiative emission that reduces the achievable open-circuit voltage $V_{\mathrm{oc}}$.[16] Multijunction photovoltaics mitigates these losses by partitioning the solar spectrum across multiple bandgaps.[1, 2, 4] In the radiative and ideal-optics limit, the efficiency of multijunction devices approaches the reciprocal multicolor limit as the number of junctions ($N$) grows.
Despite decades of progress, pushing beyond triple-junction stacks remains a challenge in conventional epitaxial material systems, because lattice matching, thermal expansion mismatch, and defect formation lead to practical constraints. Van der Waals (vdWs) semiconductors, especially transition metal dichalcogenides (TMDs), offer a fundamentally different integration pathway: their layered nature enables mechanical exfoliation, deterministic transfer printing, and assembly of many junctions without lattice matching.[3, 18, 15] This raises a natural theoretical question: if we could stack an arbitrary number of vdW junctions and could select bandgaps broadly across the solar spectrum, how close could we approach multicolor or even Landsberg limits?
Answering this question requires more than bandgap optimization. As emphasized by luminescence thermodynamics, high-efficiency photovoltaics must exhibit strong internal radiative recombination and high external photon extraction.[9, 5, 6, 14, 7] Reciprocity further implies that absorption and emission are inseparable: thickness- and nanophotonics-controlled absorptance defines not only the short-circuit current $J_{\mathrm{sc}}$ but also the radiative dark current and therefore $V_{\mathrm{oc}}$.[14, 13] In stacked junctions without intermediate mirrors,[4] an upper junction can emit externally both upward (to the environment) and downward (into the stack), which (i) opens additional optical emission channels and thus introduces an entropy and voltage penalty, and (ii) enables luminescent coupling to lower junctions.[1, 2, 4] These coupling processes can give rise to additional thermalization or entropic losses in the lower junctions, particularly when the bandgaps are closely spaced and the upper junction is highly radiative. Recent detailed-balance analyses of nanophotonic intermediate mirrors in tandem photovoltaics highlight that the net gain can be strongly device configuration-dependent, highlighting the need for stack-specific photon-entropy accounting.[4]
Motivated by this perspective, we pursue a conservative and practical goal: maximize the reciprocal multijunction efficiency achievable with a realistic TMD bandgap window (1.0–2.1 eV), and quantify how both luminescence extraction and thickness constraints define the approach to the multicolor limit. Nonreciprocal optics is treated as an optional strategy: we benchmark its theoretical efficiency gain by applying the idealized nonreciprocal multijunction model of Fan et al.[11] to the same bandgap ladders, which are optimized under reciprocity (a conservative “plug-in” comparison). Then, we systematically contrast this multijunction efficiency gain with the absence of benefit in single-junction photovoltaics.[10] These developments reinforce the motivation for a unified thermodynamic treatment of emission, coupling, and thickness constraints when designing transfer-printed TMD multijunction stacks.
2 The device concept and photon-flow constraints
Figure 1 summarizes the transfer-printed TMD multijunction architecture, studied in this work: a stack of independently contacted subcells assembled by TMD stacking or transfer printing, with selective contacts/interlayers between subcells, a transparent top electrode, and a reflective back electrode. Throughout, we focus on multi-terminal extraction (each TMD subcell operates at its own maximum power point), which isolates the fundamental spectral-partitioning thermodynamics from current-matching constraints.
Figure 1: Conceptual schematic of a transfer-printed vdW/TMD multijunction solar cell (representative $N=5$ ladder). In a reciprocal stack without intermediate mirrors, external luminescence from an upper junction can exit upward ($\Phi^{\uparrow}_{\mathrm{ext}}$) and propagate downward ($\Phi^{\downarrow}_{\mathrm{ext}}$), enabling luminescent coupling to lower junctions and opening additional emission channels (entropy/voltage penalty). Intermediate mirrors or angular filters can suppress unwanted emission channels and recover higher radiative voltage.[9, 13] An optional nonreciprocal optical layer can, in principle, further reshape emission pathways and access additional thermodynamic efficiency gain in multijunction stacks.[11]
Two points are central. First, in the detailed-balance limit, each junction is a light-emitting diode under forward bias: at open circuit and near maximum power, the device emits a substantial photon flux. Strong external luminescence is therefore not a loss to be suppressed, but a signature of a high-quality device approaching the SQ limit.[9] Second, in a stacked architecture, the direction and angular distribution of luminescence matter. If an upper junction emits downward into the stack, that emission is not necessarily “wasted.” It can be re-absorbed by lower junctions (luminescent coupling). However, it can still introduce (i) a voltage penalty for the emitter by opening emission channels, and (ii) a thermalization penalty in the absorber if the photon energy significantly exceeds the lower-junction bandgap. A quantitative framework must therefore track both photon flux and power flux of luminescence.
3 Theory and Methods
3.1 Spectral model, concentration, and incident power
We model the Sun as a blackbody at temperature $T_{s}$ and the cell as a blackbody at temperature $T_{c}$ (unless otherwise stated $T_{s}=$5778\text{\,}\mathrm{K}$\$ and $T_{c}=$300\text{\,}\mathrm{K}$\$). The spectral photon flux per unit area per unit energy per steradian is
$\Phi_{\mathrm{bb}}(E,T)=\frac{2E^{2}}{h^{3}c^{2}}\,\frac{1}{\exp(E/k_{B}T)-1}$ (1)
where $E$ is photon energy, $h$ is Planck’s constant, and $c$ is the speed of light. The Sun subtends solid angle $\Omega_{s}=\pi\sin^{2}\theta_{s}$, where $\theta_{s}\simeq$0.266\text{\,}\mathrm{d}\mathrm{e}\mathrm{g}\mathrm{r}\mathrm{e}\mathrm{e}$\$ is the solar half-angle. Optical concentration is represented by scaling the incident flux by a factor $C$, so that the spectral photon flux, which is incident on the device, is
$\Phi_{\odot}(E)=C\,\Omega_{s}\,\Phi_{\mathrm{bb}}(E,T_{s})$ (2)
The incident radiative power density is
$P_{\mathrm{in}}(C)=\int_{0}^{\infty}E\,\Phi_{\odot}(E)\,dE$ (3)
[…]
3.2 Single-junction detailed balance
For an ideal diode, the current–voltage relation is
$J(V)=J_{\mathrm{sc}}-J_{0}\left[\exp\left(\frac{qV}{k_{B}T_{c}}\right)-1\right]$ (4)
where $J_{\mathrm{sc}}$ and $J_{0}$ are the short-circuit current density and the dark-current prefactor, respectively. For an absorber with absorptance $A(E)$ (not necessarily unity), the photogenerated current density is
$J_{\mathrm{sc}}=q\int_{0}^{\infty}A(E)\,\Phi_{\odot}(E)\,dE$ (5)
Radiative recombination produces an emitted photon flux proportional to $A(E)\,\Phi_{\mathrm{bb}}(E,T_{c})$ by both Kirchhoff’s law and reciprocity.[14] Thus, the radiative dark current is
$J_{0,\mathrm{rad}}=q\int_{0}^{\infty}A(E)\,\Phi_{\mathrm{bb}}(E,T_{c})\,\Omega_{\mathrm{emit}}\,dE$ (6)
where $\Omega_{\mathrm{emit}}$ is the emission solid angle. For one-sided emission (perfect back reflector), $\Omega_{\mathrm{emit}}=2\pi$, whereas two-sided emission corresponds to $\Omega_{\mathrm{emit}}=4\pi$.
Nonradiative recombination and parasitic optical loss reduce the fraction of recombination events, which are responsible for externally emitted photons. A convenient device-quality metric is the external radiative efficiency ($\mathrm{ERE}$), which can be introduced via
$J_{0}=\frac{J_{0,\mathrm{rad}}}{\mathrm{ERE}}$ (7)
At fixed $J_{\mathrm{sc}}$, Eq. (7) produces a voltage penalty
$\Delta V_{\mathrm{ERE}}=\frac{k_{B}T_{c}}{q}\ln(\mathrm{ERE})$ (8)
which is identical in form to the $\ln Q^{\mathrm{LED}}_{e}$ term that appears in the thermodynamic voltage decomposition of Rau and co-workers.[13, 12, 6]
The maximum power point (MPP) can be obtained analytically using the Lambert-$W$ function. Starting from Eq. (4), the condition $d(VJ)/dV=0$ leads to $(1+v)\exp(v)=1+J_{\mathrm{sc}}/J_{0}$ where $v=qV/(k_{B}T_{c})$. This gives
$v_{\mathrm{mpp}}=W\!\left(e\left[1+J_{\mathrm{sc}}/J_{0}\right]\right)-1$ (9)
with $V_{\mathrm{mpp}}=(k_{B}T_{c}/q)\,v_{\mathrm{mpp}}$ and $P_{\mathrm{mpp}}=V_{\mathrm{mpp}}J(V_{\mathrm{mpp}})$. A full derivation is provided in the Supplement Information.
3.3 Unlimited-junction split-spectrum multijunction model
We consider a $N$-junction multijunction device with bandgaps $E_{g,1}>E_{g,2}>\cdots>E_{g,N}$. In the split-spectrum and multi-terminal limit, junction $i$ interacts with photon energies in a window $E\in[E_{g,i},E_{g,i-1})$, where $E_{g,0}\equiv E_{\mathrm{max}}$ is a numerical high-energy cutoff. For step absorbers, $A_{i}(E)=1$ in this window and $A_{i}(E)=0$ outside it. The total maximum power density is
$P^{\ast}_{\mathrm{tot}}=\sum_{i=1}^{N}P_{i}^{\ast}(E_{g,i};E_{g,i-1})$ (10)
where $P_{i}^{\ast}$ is obtained from Eqs. (5)–(9), which are restricted to the window. We compute optimal bandgap ladders by maximizing Eq. (10) with dynamic programming (DP) on a discretized bandgap grid (see Supplement Information). We study two scenarios: (i) unconstrained $E_{g}$ (theoretical convergence) and (ii) a conservative, experimentally available TMD bandgap window $E_{g}\in[1.0,2.1]$ eV.
3.4 Reciprocity, emission channels, and luminescent coupling in reciprocal stacks
The reciprocity relation between photovoltaic quantum efficiency and electroluminescent emission connects absorption and emission spectra.[14] A compact form for the emitted photon flux spectrum under bias $V$ is
$\phi_{\mathrm{em}}(E,V)=A(E)\,\Phi_{\mathrm{bb}}(E,T_{c})\,\Omega_{\mathrm{emit}}\left[\exp\left(\frac{qV}{k_{B}T_{c}}\right)-1\right]$ (11)
where $A(E)$ is the device absorptance (or external quantum efficiency under ideal collection). Equation (11) implies that any thickness- or nanophotonics-induced change in $A(E)$ affects both $J_{\mathrm{sc}}$ and $J_{0,\mathrm{rad}}$.
In a reciprocal multijunction stack lacking intermediate mirrors, each junction can emit both upward and downward (two hemispheres). Allowing both forward and downward hemispheres doubles the radiative dark current prefactor and produces a radiative-voltage reduction $\Delta V=(k_{B}T_{c}/q)\ln 2$ at fixed generation. More importantly, downward emission from junction $i$ can be absorbed by junction $i+1$, if $E_{g,i}>E_{g,i+1}$, producing a luminescent-coupling current $J_{\mathrm{LC}}$. We model this effect using a minimal coupled detailed-balance chain (see Supplement Information), in which
$J_{i}(V_{i})$ $=J_{\odot,i}+J^{\downarrow}_{\mathrm{rad},i-1}(V_{i-1})-\frac{J_{0,i}^{\uparrow}+J_{0,i}^{\downarrow}}{\mathrm{ERE}_{i}}\left[\exp\left(\frac{qV_{i}}{k_{B}T_{c}}\right)-1\right]$ (12)
$J^{\downarrow}_{\mathrm{rad},i}(V_{i})$ $=J_{0,i}^{\downarrow}\left[\exp\left(\frac{qV_{i}}{k_{B}T_{c}}\right)-1\right]$ (13)
with $J^{\downarrow}_{\mathrm{rad},0}\equiv 0$. This model captures (i) the emission-channel entropy penalty (i.e., two-sided emission increases $J_{0}$), and (ii) the photon recycling and coupling that can boost the lower-junction current. We solve for the system maximum power point by coordinate-ascent optimization over the set of subcell voltages $\{V_{i}\}$.
3.5 Entropy loss as upward-emitted luminescence power and coupling thermalization
A recurring conceptual challenge is to interpret luminescence in multijunction stacks: luminescence is both a necessity for high $V_{\mathrm{oc}}$ or a potential loss channel (if it can be emitted into parasitic modes). To make this quantitative, we compute the upward-emitted luminescence power
$P^{\uparrow}_{\mathrm{lum}}=\sum_{i}\int E\,\phi^{\uparrow}_{\mathrm{em},i}(E,V_{i})\,dE$ (14)
and report $P^{\uparrow}_{\mathrm{lum}}/P_{\mathrm{in}}$ as an entropy-loss proxy for emission-channel losses. In reciprocal stacks without intermediate mirrors, $P^{\uparrow}_{\mathrm{lum}}$ is nonzero, because the stack must emit externally. In the idealized nonreciprocal model, $P^{\uparrow}_{\mathrm{lum}}$ can be suppressed in multijunction architectures, enabling additional efficiency gain.[11]
Downward luminescence, in contrast, is not necessarily lost: it can be absorbed by the next junction and partially converted to electrical work. However, the portion of downward luminescent power that exceeds the electrical work extracted by the lower junction becomes heat (i.e., thermalization loss). We compute a coupling-thermalization proxy:
$P_{\mathrm{therm}}^{\mathrm{LC}}=\sum_{i=1}^{N-1}\left(P^{\downarrow}_{\mathrm{lum},i}-V_{i+1}J_{\mathrm{LC},i}\right)$ (15)
where $P^{\downarrow}_{\mathrm{lum},i}$ is the downward luminescent power emitted by junction $i$, while $J_{\mathrm{LC},i}$ is the corresponding coupling current absorbed by junction $i+1$. We use these definitions to separate upward entropy loss from coupling-induced thermalization for a representative $N=5$ ladder.
3.6 Finite absorptance, excitonic TMD absorption, and thickness constraints
Although the SQ limit assumes unity absorptance above $E_{g}$, real TMD absorbers are ultrathin (less than 1 nm for a monolayer); consequently, their absorption cannot be unity. Also, it is dominated by strong excitonic resonances and large absorption coefficients in the visible; nevertheless, tens-of-nanometers thickness is typically required to approach unity absorptance over broad spectral windows without strong light trapping.[15, 3] We therefore supplement the step-absorber model with a thickness-dependent absorptance model:
$A(E;t)=1-\exp[-\alpha(E)L_{\mathrm{eff}}(E)],\quad L_{\mathrm{eff}}(E)=F(E)\,t$ (16)
where $\alpha(E)$ and $F(E)$ are the absorption coefficient and an optional path-length enhancement factor, respectively. To capture TMD-like spectral shape, we adopt a minimal excitonic absorption coefficient model consisting of an Urbach tail, a continuum onset, and A/B exciton peaks (see Supplement Information). Because reciprocity requires the same $A(E;t)$ in Eq. (6), thickness directly affects both $J_{\mathrm{sc}}$ and $V_{\mathrm{oc}}$.
In the ultrathin regime, geometric-optics light-trapping factors (e.g., $4n^{2}$, where $n$ is the refractive index) may be overly optimistic. At an elementary level, the path-length enhancement factor is simply $F(E)=L_{\mathrm{eff}}(E)/t$, i.e., the ratio between the average optical distance traveled by photons inside the absorber and the physical thickness $t$: $F(E)=1$ describes a single-pass planar film, while $F(E)>1$ reflects multi-pass propagation and light trapping (back reflectors, resonant cavities, and scattering/angle randomization). In thick absorbers with randomized ray directions, the broadband average is bounded by the Yablonovitch limit $F\leq 4n^{2}$.[17]
When the thickness becomes comparable to the wavelength, however, light is better viewed as coupling through a finite set of radiative and guided modes, and the number of usable optical channels grows with thickness. Yu, Raman, and Fan derived wave-optics bounds on the maximum achievable enhancement in nanophotonic textures as a function of thickness and in-plane periodicity,[17] and Miller emphasized the more general thickness–functionality connection for broadband optics.[8] To capture this thickness-limited behavior without committing to a specific nanophotonic design, we introduce a normalized broadband enhancement function $g(t)\in[0,1]$ and define a thickness-dependent path-length enhancement factor:
$F_{\mathrm{proxy}}(t)=1+\left(4n^{2}-1\right)g(t)$ (17)
so that $F_{\mathrm{proxy}}=1$ in the no-light-trapping limit and $F_{\mathrm{proxy}}\to 4n^{2}$ only when sufficient thickness (and optical degrees of freedom) is available. In this work, we use $n=4.5$ as a representative TMD $n$ value (in general, $n=4.5$ of representative TMDs are ranged from 4.17 to 4.70, as summarized in Table S2, Supplement Information) and take a simple saturating proxy $g(t)=1-\exp(-t/t_{0})$ with $t_{0}=200~\mathrm{nm}$ (see Supplement Information). When using this proxy in Eq. (16), we approximate $F(E)\!\approx\!F_{\mathrm{proxy}}(t)$, and also discuss how wavelength-resolved bounds can be incorporated, when a specific device architecture is known (see Supplement Information).
4 Results
Efficiency versus junction $N$ and the role of the bandgap window. Figure 2 summarizes the detailed-balance efficiency versus junction number $N$ for (i) unconstrained bandgaps and (ii) the conservative TMD bandgap window (1.0–2.1 eV), under both 1 sun and full concentration. Under full concentration, the unconstrained efficiency increases from $\sim 40\%$ (single junction) to $\sim 70.7\%$ for $N=5$ and $\sim 84.5\%$ by $N=50$ (approaching the reciprocal multicolor limit). Imposing the TMD window yields a pronounced large-$N$ plateau: $\sim 61.5\%$ for $N=5$ and $\sim 63.4\%$ by $N=50$. This plateau reflects an intrinsic materials constraint: once the accessible bandgap set no longer spans the entire solar spectrum, additional junctions provide diminishing returns.
Physically, the reciprocal multicolor limit is reached, when the spectrum assigned to each junction becomes narrower as $N$ increases. This multijunction approach suppresses thermalization, while maintaining a high radiative voltage under concentration. In the conservative TMD window, photons below 1.0 eV (near-IR) and above 2.1 eV (blue/UV) are irreversibly lost to transmission, so increasing $N$ primarily repartitions the same limited spectral band. The rapid saturation by $N$ on the order of five, therefore, indicates that, within TMD-only bandgaps, further progress hinges more on luminescence quality (ERE), emission-channel control, and ultrathin optics than on pushing junction count. For example, under full concentration, the TMD-window ladder rises from $\eta\approx 54.2\%$ at $N=2$ to $\eta\approx 61.5\%$ at $N=5$, whereas increasing junction $N$ to $10$ enables only $\eta\approx 63.4\%$ (Table 1).
Figure 2: Detailed-balance efficiency versus junction number. (a) 1 sun and (b) full concentration for unconstrained optimal bandgaps and a conservative TMD bandgap window (1.0–2.1 eV), computed for split-spectrum (multi-terminal) multijunctions with one-sided emission and $\mathrm{ERE}=1$. Numerical tables and bandgap ladders are provided in the SI and the accompanying code package.
Efficiency limit versus accessible bandgap window. To quantify how strongly the accessible bandgap window constrains large-$N$ performance, Figure 3 maps the efficiency limit (herein, shown for $N=20$ at full concentration) versus $(E_{g,\min},E_{g,\max})$. The conservative TMD bandgap point (1.0–2.1 eV) are located far from an unconstrained optimum, explaining the plateau in Figure 2. This map provides a compact way to evaluate how expanding the accessible bandgap range (e.g., by including other vdW semiconductors beyond conventional TMDs, or by alloying/strain) could raise the efficiency plateau.
The map also highlights an asymmetry between extending the lowest- and highest-bandgap bounds. Lowering $E_{g,\min}$ generally leads to a larger efficiency gain, because the solar photon flux is relatively high in the near-infrared (NIR), whereas raising $E_{g,\max}$ mainly improves voltage and reduces high-energy thermalization. This suggests that identifying a stable narrow-gap bottom absorber in the 0.8–1.0 eV range can provide disproportionate leverage, even if the top bandgap remains near 2 eV. Physically, the lower cutoff $E_{g,\min}$ directly sets the fraction of NIR photons that are transmitted without conversion; within a fixed-width window, extending $E_{g,\min}$ downward, therefore, provides a larger efficiency gain potential than modestly increasing $E_{g,\max}$.
Figure 3: Efficiency limit versus accessible bandgap window. Heatmap of the maximum achievable efficiency for a representative high-NN case ($N=20$, multi-terminal, $\mathrm{ERE}=1$, full concentration), as a function of the allowed bandgap window $(E_{g,\min},E_{g,\max})$. The conservative TMD window (1.0–2.1 eV) is highlighted.
Optimal bandgap ladders and a representative $N=5$ TMD design. Figure 4 shows optimal bandgap ladders for selected $N$ within the conservative TMD window. As $N$ increases, the ladder becomes denser, with the top junction pushed toward the maximum available gap ($\sim 2.1$ eV) and the bottom junction pushed toward $\sim 1$ eV.
We adopt $N=5$ as a representative experimentally achievable design (transfer-printed stacks beyond triple junction are a key motivation of this work). For $N=5$ in the TMD window, the DP optimizer yields the ladder
$E_{g}^{(N=5)}\approx(2.10,\,1.78,\,1.50,\,1.24,\,and\ 1.00)\ \mathrm{eV}\quad\text{(from top to bottom)}$ (18)
This ladder is used throughout subsequent analyses of luminescence, coupling, and thickness. Layer-resolved operating metrics for this representative $N=5$ stack (including ERE and finite-absorptance effects) are provided in the Supplement Information (Figures S1 and S2).
For the representative $N=5$ ladder, the rung spacing is not uniform. The optimizer concentrates rungs, where the solar spectrum has relatively high photon flux, while keeping a comparatively wide spacing between the lowest two bandgaps. This nonuniformity matters beyond spectral splitting: in reciprocal stacks without intermediate mirrors, smaller bandgap spacing increases downward luminescent coupling and the associated thermalization burden in the lower cell. Thus, optical design choices that control two-sided emission can be as important as the bandgap selection itself.
Figure 4: Optimal bandgap ladders in the conservative TMD window. $E_{g,i}$ vs subcell index for selected $N$ values (top cell is index 1). The $N=5$ ladder in Eq. (18) is highlighted as a representative design.
Material mapping: candidate vdW and TMD absorbers for the $N=5$ ladder. To connect target bandgaps to realizable absorbers, Figure 5 overlays representative optical bandgaps, which were reported for a selection of vdW semiconductors and TMD thickness variants.[3, 15] This mapping is necessarily approximate, because optical bandgaps depend on thickness, strain, dielectric environment, and exciton binding energy. Nevertheless, this comparison still provides a practical starting point for selecting TMD subcell candidates for stacks.
Two practical subtleties are noteworthy, when translating target optical bandgaps into device layers. First, monolayer TMD optical bandgaps are excitonic and depend on dielectric environment and strain, so the same nominal absorber can shift in both absorptance and radiative efficiency, depending on encapsulation and adjacent layers. Second, many multilayer TMDs tend to retain indirect-bandgap as thickness increases, which can reduce luminescence and therefore reduce voltage, unless light management and passivation mitigate nonradiative recombination. Among the five rungs, the bottom cell target near 1.0 eV is the most challenging within conventional TMDs and is also the most consequential for the efficiency limit, because it controls NIR photon harvesting. This makes contact engineering, encapsulation, and selection of narrow-gap vdW absorbers central to realizing the theoretical efficiency gain.[3, 15, 18] Table 1 provides a compact rung-to-material mapping for the representative $N=5$ ladder, emphasizing TMD-first choices within the conservative 1.0–2.1 eV window. The broader candidate database, used in the thickness-dependent calculations, such as thickness, strain, alloy tunability ranges, and optical-constants, is provided in the Supplement Information (Tables S3–S6).
Figure 5: Illustrative mapping between target bandgaps and vdW/TMD candidates. Symbols denote representative optical gaps drawn from the vdW/TMD literature (see Supplement Information, Table S3 for notes and sources), and dashed vertical lines denote the representative $N=5$ target ladder (Eq. (18)).
Table 1: Representative rung-to-material mapping for the $N=5$ target ladder within the conservative 1.0–2.1 eV window. Primary choices emphasize widely studied group-VI TMDs compatible with transfer printing; alternates illustrate thickness/strain/alloy routes and broader vdW options. See the Supplement Information, Tables S3–S6 for a more comprehensive candidate library and parameter notes.
Target $E_{g}$ (eV) Rung Primary candidate Alternates Practical notes (tunability / caveats)
2.10 Top WS2 (1L) MoS2 (1L, strain/dielectric); HfS2 (few-L) Wide-gap top cell; strong visible excitons; encapsulation and work-function control mitigate contact loss.
1.78 2 MoS2 (1L) WS2 (2L–few-L); MoS2(1-x)Se2x (alloy) Gap tunable by thickness and dielectric screening; prioritize high ERE and low series resistance.
1.50 3 MoSe2 (1L) ReS2; WSe2 (1L, strain) Mid-gap cell with strong excitonic absorption; minimize parasitic absorption in interlayers.
1.24 4 WSe2 (few-L/bulk) MoS2 (bulk); InSe (few-L) Near-IR edge; multilayer TMDs may become indirect; thickness increase can compensate finite absorptance.
1.00 Bottom MoTe2 (few-L/bulk) MoSe2 (bulk, $\sim$1.1); BP (encapsulated, thickness-tuned) Narrow-gap bottom cell; stability/oxidation and contact barriers require encapsulation and barrier-free contacts.
System-level penalty from finite ERE. Figure 6 shows how efficiency degrades as $\mathrm{ERE}$ decreases below unity, for selected $N$ values in the conservative TMD window. Because the voltage penalty is logarithmic, efficiencies remain relatively robust for $\mathrm{ERE}\gtrsim 10^{-2}$, but decline rapidly for lower $\mathrm{ERE}$. Importantly, in multijunction systems, the penalty compounds across junctions: one low-ERE subcell can dominate the system loss.[9, 12, 6] This motivates aggressive light-management and contact designs (e.g., low parasitic absorption, high-reflectivity mirrors, and low-loss transparent electrodes) that preserve external photon extraction in every TMD subcell.
A simple quantitative rule-of-thumb illustrates the compounding nature of the loss: at room temperature, an ERE of 1e-2 corresponds to an open-circuit voltage penalty on the order of one tenth of a volt per junction, so a five-junction stack can lose several tenths of a volt in total, even if current matching is ideal. This is why maintaining high ERE in every TMD subcell (not only the top) is a prerequisite for approaching multicolor efficiencies. For the representative $N=5$ ladder, reducing ERE from 1 to $10^{-2}$ lowers the full-concentration stack efficiency by about six absolute percentage points in the thick-limit regime (Figure 7).
Figure 6: Efficiency versus external radiative efficiency (ERE) in the TMD window. Efficiency computed for $N=1,2,5$ under 1 sun (solid) and full concentration (dashed), illustrating the SQ-triangle voltage penalty and its correspondence to the thermodynamic $Q^{\mathrm{LED}}_{e}$ term.[13, 12]
Thickness dependence with excitonic TMD absorptance. Figure 7 moves beyond step absorbers to incorporate finite absorptance in ultrathin TMD layers. Panel (a) shows the excitonic absorption coefficient model used herein, including A/B exciton peaks and a continuum onset; the model is intentionally minimal but captures the qualitative spectral features of many group-VI TMDs.[3, 15] The underlying optical-constant conventions ($n$, $k$, and $\alpha$) and representative material parameters used in this work are summarized in the Supplement Information (Tables S2–S6). Panel (b) shows the resulting $N=5$ stack efficiency versus a uniform TMD subcell thickness $t$ for $\mathrm{ERE}=1$ and $\mathrm{ERE}=10^{-2}$. Even though monolayer and few-layer TMDs exhibit pronounced excitonic absorption near the band edge, the usable solar spectrum spans a broad range of above-bandgap energies and the continuum absorption remains finite; consequently, a very thin film does not absorb all incident photons in a single pass. In the Beer–Lambert regime ($\alpha t\!\ll\!1$), the absorptance scales approximately as $A\approx\alpha t$, so the short-circuit current and efficiency rise only gradually with thickness.
Using a 95% convergence criterion relative to the thick-absorber (step) limit for the full-concentration $N=5$ ladder, the required thickness is $\sim 120$ nm per TMD subcell for a single-pass geometry ($F=1$). When broadband light trapping is included through the thickness-dependent proxy of Eq. (17) (i.e., $F(E)\approx F_{\mathrm{proxy}}(t)>1$ constructed from the normalized enhancement $g(t)$ in Figure 8c), this target thickness drops to $\sim 20$ nm. This shift provides an experimentally important design message: moderate broadband path-length enhancement can trade ”optical” thickness for ”physical” thickness in ultrathin TMD stacks. Finite absorptance imposes a double penalty in detailed balance. First, incomplete absorption directly reduces $J_{\mathrm{sc}}$ by allowing above-bandgap photons to transmit. Second, because the same absorptance weights radiative emission channels, reduced absorption increases the effective recombination-to-generation ratio that sets the voltage (the $J_{\mathrm{sc}}/J_{0}$ ratio) and lowers both $V_{\mathrm{mpp}}$ and the overall power. The proxy $F_{\mathrm{proxy}}(t)$ accelerates convergence by increasing the effective photon dwell time, but it cannot make an arbitrarily thin absorber behave as a thick one: the normalized enhancement $g(t)$ is itself thickness-limited by nanophotonic channel counting and thickness–functionality bounds (Figure 8).[17, 8, 5]
Figure 7: Thickness dependence for an $N=5$ TMD multijunction stack with excitonic absorptance. (a) Example absorption coefficient model $\alpha(E)$ for a representative bandgap ($E_{g}=1.78$ eV) including A/B exciton peaks and a continuum onset. (b) Total $N=5$ stack efficiency under full concentration versus uniform subcell thickness $t$, comparing single-pass absorption ($F=1$) to the thickness-dependent broadband light-trapping proxy $F_{\mathrm{proxy}}(t)$ of Eq. (17). Reciprocity is enforced by using the same $A(E;t)$ in both $J_{\mathrm{sc}}$ and the radiative dark current.
Nanophotonic bounds and the role of thickness. The need for a thickness-dependent broadband proxy follows from the distinction between geometric optics and wavelength-scale optics. In thick absorbers with randomized ray directions, the average internal path length can be increased up to the Yablonovitch bound $F\leq 4n^{2}$. When $t\lesssim\lambda$, however, the optical response is mediated by a finite set of radiative and guided modes; the number of accessible channels (and therefore the achievable broadband response) grows with thickness. As a result, nanophotonic textures may yield strong enhancement at selected wavelengths or angles, but sustaining large enhancement over the broad solar spectrum is fundamentally thickness-limited.
Figure 8 summarizes representative results from nanophotonic light-trapping theory and motivates the normalized enhancement function $g(t)$ used throughout this work. Figure 8(a) illustrates the Yu–Raman–Fan wave-optics enhancement oscillations around the $4n^{2}$ limit for periodic textures, highlighting the discrete-channel nature of wavelength-scale light trapping.[17] Figure 8(b) illustrates single-mode scaling trends emphasized in nanophotonic bounds.[17] Figure 8(c) plots the simple saturating function $g(t)$ adopted here, which can be interpreted as the fraction of the $4n^{2}$ enhancement that is experimentally achievable as a broadband average at thickness $t$. We then construct the effective path-length enhancement factor $F_{\mathrm{proxy}}(t)=1+(4n^{2}-1)g(t)$ [Eq. (17)], which is used in the absorptance model of Eq. (16) and in the thickness sweeps of Figure 7.[8, 5] For multijunction stacks, this thickness constraint is wavelength selective: the longest-wavelength TMD subcells are the hardest to enhance in the ultrathin regime, so the bottom junction often dictates the minimum practical thickness or the required strength of light trapping. As such, optical design should be co-optimized with the bandgap ladder, rather than treated as a post-processing step.[8]
Figure 8: Nanophotonic bounds and a broadband path-length enhancement proxy. (a) Channel-counting proxy illustrating deviations from the geometric-optics $4n^{2}$ limit in the wavelength-scale regime. (b) Single-mode scaling proxy for $F_{\max}/(4n^{2})$ versus normalized thickness $\tau=d/(\lambda/2n)$. (c) Normalized broadband enhancement $g(t)=(F_{\mathrm{proxy}}(t)-1)/(4n^{2}-1)$ used to construct the thickness-dependent path-length enhancement factor $F_{\mathrm{proxy}}(t)=1+(4n^{2}-1)g(t)$ in Figure 7.[17, 8, 5]
Luminescent coupling and a power-budget view of entropy vs thermalization. Figure 9 quantifies when luminescent coupling becomes appreciable in reciprocal stacks and provides a power-budget decomposition for the representative $N=5$ ladder. Figure 9(a) shows that the downward coupling current $J_{\mathrm{LC}}$ at the system MPP increases strongly with concentration and is amplified at high ERE due to higher operating voltage. Figure 9(b) shows that the coupling ratio grows rapidly, when adjacent bandgaps become closely spaced (small $\Delta E_{g}$) (i.e., a condition that is naturally approached in dense high-$N$ ladders). Figure 9(c) provides a compact power budget for the representative $N=5$ ladder: the upward luminescence power is plotted as an entropy-loss proxy, while coupling-induced thermalization is plotted separately. This decomposition clarifies the key design point: suppressing upward emission can reduce entropy loss (and increase voltage), but downward emission into lower junctions can still lead to thermalization loss, if energy is not efficiently converted.
Importantly, coupling is not universally detrimental. When the lower cell converts the coupled photons efficiently and the bandgap spacing is sufficiently large, downward luminescence can partially compensate current deficits and relax the need for perfect spectral splitting. However, in dense ladders with small bandgap spacing, coupled photons arrive with excess energy relative to the lower bandgap. As a result, these coupled photons are more likely to be dissipated as heat, effectively reintroducing a thermalization channel that grows with concentration and ERE. For the representative $N=5$ ladder at full concentration, the reciprocal ideal-optics stack loses $\sim 1.24\%$ of the incident solar power as upward-emitted luminescence, whereas in the idealized nonreciprocal model, this upward channel is eliminated and the electrical output increases by $\sim 1.5\%$ absolute, with part of the redirected luminescence, which appear as coupling-induced thermalization (Figure 9c).
Figure 9: Luminescent coupling conditions and entropy/thermalization power budget. (a) Downward luminescent coupling current at MPP increases strongly with optical concentration and is larger for higher ERE. (b) Coupling ratio $\gamma=J_{\mathrm{LC}}/J_{\odot,2}$ increases rapidly as the bandgap spacing decreases. (c) Power-budget decomposition for the representative $N=5$ ladder under full concentration, comparing reciprocal ideal optics (no coupling), reciprocal stacks without intermediate mirrors (two-sided emission + coupling), and an idealized nonreciprocal model. Upward-emitted luminescence is shown as an entropy-loss proxy; coupling thermalization quantifies luminescent power not convertible to electrical work in the lower junctions.
Reciprocal vs nonreciprocal efficiency gain. Nonreciprocal multijunction photovoltaics has been proposed as a route to approach the Landsberg limit by suppressing entropy losses, which are associated with reciprocal emission pathways.[11] However, for single-junction solar cells, nonreciprocity does not increase the SQ efficiency limit under detailed balance.[10] To place these insights in the present context, we benchmark the potential efficiency gain by applying the idealized nonreciprocal multijunction model to the same bandgap ladders, that are optimized under reciprocity (Option A, conservative comparison).
Figure 10 compares full-concentration efficiencies versus $N$ for reciprocal and nonreciprocal cases, for both unconstrained bandgaps and the conservative TMD window. In the TMD window, the reciprocal efficiency saturates near $\sim 63.4\%$ at large $N$, whereas the nonreciprocal model reaches to $\sim 67.7\%$ at $N=50$ (Figure 10a): extended bandgap ladder-tables are available in Tables S7–S10, Supplement Information. Figure 10b shows the critical physical signature: the nonreciprocal model suppresses upward-emitted luminescence power (entropy-loss proxy), whereas the reciprocal case must emit upward to maintain detailed balance. The filled markers in Figure 10a further show that reciprocal stacks without intermediate mirrors incur additional penalties from two-sided emission and coupling, highlighting the practical importance of intermediate mirrors or angular filters[4] in realistic reciprocal designs. In the conservative Option A comparison, the nonreciprocal model is implemented to the reciprocity-optimized ladder without re-optimizing bandgaps, so the improvement should be interpreted as a lower bound on the possible efficiency gain. In practice, achieving the ideal nonreciprocal behavior is nontrivial, but the comparison clarifies what must be suppressed in reciprocal stacks: upward emission that carries entropy and downward emission that drives coupling thermalization. Intermediate mirrors, angular filters, and spectral-selective contacts[4] can address part of this gap without requiring fully nonreciprocal photonics.
Figure 10: Reciprocal vs nonreciprocal efficiency gain under full concentration. (a) Efficiency versus junction number for reciprocal split-spectrum multijunctions (one-sided emission, no coupling), and for the idealized nonreciprocal multijunction model of Fan et al.[11] applied to the same reciprocity-optimized bandgap ladders (Option A). Filled markers show a reciprocal-stack estimate without intermediate mirrors (two-sided emission + downward luminescent coupling) for $N\leq 10$. (b) Upward-emitted luminescence power fraction $P^{\uparrow}_{\mathrm{lum}}/P_{\mathrm{in}}$ as an entropy-loss proxy. In the idealized nonreciprocal model used here, the upward-emitted luminescence is identically zero, so the nonreciprocal series in (b) are shown with a slight horizontal offset for visibility. Horizontal lines indicate literature multicolor and Landsberg limits for comparison (values depend weakly on $T_{s}$).[11]
5 Discussion: design rules for TMD multijunction cells
(1) Bandgap-ladder engineering is necessary but not sufficient.
The DP optimizer provides optimal ladders and reveals the large-$N$ plateau imposed by a constrained bandgap window. Within the conservative 1.0–2.1 eV TMD window, $N\approx 5$ already approaches the efficiency plateau, suggesting that the most immediate experimental objective is to realize high-quality $N\leq 5$ stacks with excellent optics rather than extremely large $N$.
(2) ERE is a system-level requirement.
Voltage penalties from finite ERE accumulate across subcells (Figure 6). High-NN stacks, therefore, demand high radiative quality in every junction, highlighting the importance of passivation, low-loss contacts, and photon extraction.[9, 6]
(3) Control of emission channels and luminescent coupling are essential.
Without intermediate mirrors, two-sided emission introduces an entropy penalty and enables downward luminescent coupling (Figure 9a–b). Coupling can be beneficial for current in lower cells. However, it also introduces thermalization losses and reduces the voltage. Practical multijunction architectures will likely require selective optical elements (intermediate mirrors, angular filters, or photonic cavities) designed to maximize useful coupling while minimizing entropy penalties.[13, 17, 4]
(4) Thickness and nanophotonics cannot be ignored in vdW stacks.
Ultrathin absorbers cannot be treated as step absorbers; finite absorptance reduces both $J_{\mathrm{sc}}$ and $V_{\mathrm{oc}}$ through reciprocity (Figure 7). Nanophotonic bounds and thickness constraints restrict the achievable broadband path-length enhancement (Figure 8). These effects motivate moderate absorber thicknesses (tens to hundreds of nm) and co-designed photonic structures rather than extreme monolayer limits for high-efficiency multijunction photovoltaics.
(5) Nonreciprocity is an optional asymptotic strategy with multijunction-specific benefits.
Nonreciprocity does not improve the SQ limit for single-junction devices.[10] In multijunction architectures, however, rerouting emission pathways can reduce entropy losses and provide efficiency gain toward Landsberg-like limits.[11] Our conservative plug-in comparison (Figure 10) quantifies this efficiency gain for TMD-optimized ladders, while preserving our work’s main scope: reciprocal multijunction design and realistic constraints.
6 Conclusions
We developed an unlimited-junction detailed-balance framework tailored to transfer-printed TMD multijunction photovoltaics and used it to compute optimal bandgap ladders under unconstrained and conservative TMD bandgap windows. A realistic bandgap window produces a large-$N$ efficiency plateau, motivating $N\sim 5$ as a practical near-term target for TMD-based stacks. We incorporated luminescence thermodynamics and reciprocity to quantify ERE penalties, emission-channel entropy, and luminescent coupling. Also, we introduced a power-budget view, in which upward-emitted luminescence serves as a directly computable entropy-loss proxy, while coupling thermalization is tracked, separately. Finally, we benchmarked the remaining thermodynamic efficiency gain by applying an idealized nonreciprocal multijunction model to our reciprocally optimized ladders, showing negligible benefit for single junctions but substantial efficiency gain for multijunction stacks. These results provide quantitative design rules for approaching multicolor efficiencies with transfer-printed TMD multijunction photovoltaics.
Data and code availability
The numerical results and figures in this work are reproducible from the Python implementation of the detailed-balance, luminescent-coupling, and optimization models described in the main text and Supplement Information. The code is available at https://github.com/NEOlab-code/Multijunction-TMD-solar-cells.git. To keep the manuscript self-contained for submission, the Supplement Information includes the full derivations, parameter tables, and additional figures. The Python source code and the raw numerical data required to regenerate every figure and table are available from the authors upon reasonable request and will be deposited in a public repository upon publication.
Supplement Information
The Supplement Information is provided as a separate PDF file. The SI provides detailed derivations, parameter tables, and additional figures that support the main text:(i) Section 1 (constants, geometry, and spectral model): solar/blackbody spectral definitions and constants; (ii) Section 2 (single-junction detailed balance): full detailed-balance derivations, including the Lambert-$W$ maximum-power condition; (iii) Section 3 (unlimited-junction split-spectrum multijunction model): the dynamic-programming (DP) optimizer used to compute optimal bandgap ladders for arbitrary $N$ and its discretization/convergence checks; (iv) Section 4 (reciprocity, emission channels, and the role of mirrors): reciprocity-based voltage-loss and entropy-loss decompositions, including the role of external radiative efficiency (ERE) and optical outcoupling; (v) Section 5 (luminescent coupling chain model): the luminescent-coupling chain model, together with the definition and computation of the upward emitted luminescence power and a coupling thermalization/heat proxy; (vi) Section 6 (TMD optical response and excitonic absorptance model): a TMD material database (representative bandgaps, excitonic absorption spectra, and representative $n(\lambda)$$k(\lambda)$ dispersion conventions) and mapping of candidate stacks to the representative $N=5$ target ladder; (vii) Section 7 (nanophotonic bounds and Miller thickness constraints): nanophotonic bounds and thickness constraints relevant to ultrathin absorbers; (viii) Section 8 (nonreciprocal multijunction benchmark); (ix) Section 9 (extended bandgap-ladder tables for reproducibility); (x) Section 10 (additional figures for the representative $N=5$ stack; Figures S1 and S2); (xi) Section 11 (full ladder lists in text form).
Author Information
Corresponding Author
Seungwoo Lee — E-mail: seungwoo@korea.ac.kr
Author Contributions
S.L. conceived the original research idea at the initial stage, computed all the results, and wrote the paper.
Notes
The author declares no competing financial interest.
Acknowledgments
We acknowledge funding from National Research Foundation of Korea (RS-2022-NR068141) and from the KIST Institutional Program (Project No.: 2V09840-23-P023). This research was also supported by a grant of the Korea-US Collaborative Research Fund (KUCRF), funded by the Ministry of Science and ICT and Ministry of Health & Welfare, Republic of Korea (grant number: RS-2024-00468463) and by Korea University grant.